if the length of a rectangle is 4 feet and the perimeter is 30 feet what is the area?
Answer by checkley77(12844)

L=4 FT.
30=2L+2W
30=2*4+2W
30=8+2W
2W=30-8
2W=22
W=22/2
W=11 FT. ANS. FOR THE WIDTH.
AREA=LW
AREA=4*11=44 FT^2. ANS. FOR THE AREA.